OpenStudy (iiariana):

What is the volume of the box if it is scaled down by a factor of 1/10? BOX VOLUME 320 in3

Vocaloid (vocaloid):

this is a bit of a tricky question when we scale down by a factor of 1/10, we mean that we multiply each dimension by (1/10) old volume = L*W*H new volume = (1/10)L*(1/10)W*(1/10)*H = (1/1000)L*W*H new volume = (1/1000)*old volume